1) the equation of a curve is given as f(x)=√2x-1.find the equation of the tangent line to the graph of f(x) at x=5.


2) find the equation of the tangent line to the curve g(x)=x/x+1 at the point (2,2/3).


1
Expert's answer
2021-05-07T14:13:58-0400

(1) Given curve is-

"y=\sqrt{2x-1}"

"\Rightarrow y'(x)=\dfrac{2}{2\sqrt{2x-1}}\\[9pt]\Rightarrow y'(x)=\dfrac{1}{\sqrt{2(5)-1}}=\dfrac{1}{\sqrt{9}}=\dfrac{1}{3}"


Also, at "x=5, y=\sqrt{2(5)-1}=\sqrt{10-1}=\sqrt{9}=3"


So Tangent Equation is-


"(y-3)=y'(x)(x-5)\\[9pt]

\Rightarrow y-3=\dfrac{x-5}{3}\\[9pt]\Rightarrow 3y-9=x-5\\[9pt]\Rightarrow 3y-x-4=0"


(2) Given curve is-


"g(x)=\dfrac{x}{x+1}"


"g'(x)=\dfrac{(x+1)-x}{(x+1)^2}\\[9pt]\Rightarrow g'(x)_{(2,\frac{2}{3})}=\dfrac{1}{(2+1)^2}=\dfrac{1}{9}"


Tangent Equation is-


"(y-\dfrac{2}{3})=g'(x)(x-2)\\[9pt]\Rightarrow \dfrac{3y-2}{3}=\dfrac{x-2}{9}\\[9pt]\Rightarrow 9y-6=x-2\\[9pt]\Rightarrow x-9y+4=0"